Circular No. 8748

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET P/2006 R2 (CHRISTENSEN) E. J. Christensen reports his discovery of a comet on Catalina Sky

Survey images taken with the 0.68-m Schmidt telescope, on which the new object displays a 15′′ coma with a short tail toward p.a. 300^{o} in four 30-s stacked images (discovery observation tabulated below). Following posting on the ‘NEO Confirmation Page’, several other observers have noted the object’s cometary appearance on their CCD images, including R. H. Mc-Naught (Siding Spring, 0.5-m Uppsala Schmidt telescope; Sept. 14.57 UT, 12” coma with 20” tail in p.a. 290^{o}), C. Jacques and E. Pimentel (Belo Horizonte, Brazil, 0.30-m f/3 Schmidt-Cassegrain telescope; Sept. 15.14– 15.16, central condensation in a coma of diameter 22′′ and a small faint tail 50′′ long in p.a. 280^{o}), and J. E. McGaha (Tucson, AZ, 0.36-m f/10 Schmidt-Cassegrain reflector; Sept. 15.4, faint, round 4′′ coma with a faint 8′′ elongation in p.a. 300^{o} in ten stacked 60-s images).

2006 UT  α_{2000}  δ_{2000}  Mag.
Sept.14.32179  23^{h}16^{m}44.^{s}03  −16^{o}45′49′′.6  17.5

The available astrometry, the following preliminary orbital elements, and an ephemeris appear on MPEC 2006-R52.

T = 2006 June 30.430 TT  ω = 192.^{o}397
e = 0.27891  Ω = 138.283 }2000.0
q = 2.98538 AU  i = 15.688
a = 4.14011 AU n^{o} = 0.117000 P = 8.4 years

SW URSAE MAJORIS E. O. Waagen, AAVSO, reports that the cataclysmic variable SW

UMa is in superoutburst for the first time since Oct. 2002, when the star reached visual mag 10.5 and was brighter than mag 13.5 for 17 days (the normal minimum visual mag being ≈ 17). Visual magnitudes reported to the AAVSO: Aug. 21.876, [13.5 (D. Naillon, Lithons, France); Sept. 12.140, [13.0 (P. Schmeer, Bischmisheim, Germany); 13.149, 10.8 (Schmeer); 14.036, 10.2 (S. Swierczynski, Dobczyce, Poland); 14.138, 10.4 (Schmeer); 14.172, 10.4 (Schmeer); 15.024, 10.1 (Swierczynski). Waagen adds that the outburst frequency of SW UMa varies a lot, from less than a year to as long as 5 years, as does the pattern of normal outbursts versus superoutbursts; prior to the 2002 outburst, there had been an outburst only 15 months previously.

2006 September 15  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2006 CBAT  Daniel W. E. Green